# section[1.5em]1.5em1pc

### v2

@@ 10.2.   Operadic point of view
We now use the characterisation of $\Gamma(\operatorname{Lie})$-algebras using 3.3. Following [Ikonicoff23]*Example 6.6.c), if $L$ is a $\Gamma(\operatorname{Lie})$-algebra, the Lie bracket on $L$ is given by $\beta_{[-,-],(1,1)}$, and the $p$-map is given by $\beta_{F_{p},(p)}$, where $F_{p}\in\operatorname{Lie}(p)^{\Sigma_{p}}$ is the element:
@@
where $\sigma$ ranges over $\Sigma_{1}\times\Sigma_{p}$ in the sum, which can also be written $\beta_{[-,-],(1,1)}(\beta_{1_{\mathcal{P}},(1)},\beta_{F,(p)})=\beta_{F_{p+1},(1,p)}$.
@@
Following Definition 4.5, an $L$-module $M$ becomes a vector space $M$ equipped with a semilinear map $f\colon M\to M$ and an action of $L$ that we denote $[l,m]$ for $l\in L$, $m\in M$ such that $[l,f(m)]=0$. This last relation comes from the following computation:
$[l,m^{[p]}]$ $=\beta_{[-,-],(1,1)}(l,\beta_{F,(p)}(m)),$
$=\beta_{\sum_{\sigma}\sigma[-,-]\circ_{1}\cdots\circ_{1}[-,-],(1,p)}(l,m),$
@@
Note that $L$ injects into $\mathbb{U}_{\Gamma(\operatorname{Lie})}(L)$, by identifying $l\in L$ to $\beta_{[-,-],(1,1)}(l,-)$. For an $L$-module (i.e., a $\mathbb{U}_{\Gamma(\operatorname{Lie})}(L)$-module) $M$, we will denote by $[l,m]:=\beta_{[-,-],(1,1)}(l,m)=\beta_{[-,-],(1,1)}(l,-)\cdot m$, and $f(m):=\beta_{F_{p},(p)}(m)=\beta_{F_{p},(p)}(-)\cdot m$ for $l\in L$, $m\in M$.
@@
Following Proposition 7.2, a Beck derivation $d\colon L\to M$ is a linear map $d\colon L\to M$ satisfying $d\left([l,l^{\prime}]\right)=\beta_{[-,-],(1,1)}(l,dl^{\prime})+\beta_{[-,-],(1,1)}(dl,l^{\prime})=[l,dl^{\prime}]-[l^{\prime},dl]$, and
$d\left(l^{[p]}\right)=\sum_{i=1}^{p}\beta_{F_{p},(p-i,i)}(l,dl).$
Note that, since $F_{p}\in Lie(p)^{\Sigma_{p}}$, for all $i$ such that $12\leq i\leq p-$ 1,
@@
and according to relation ($\beta$M4), $\beta_{\sum_{\tau\in\Sigma_{p-i}\times\Sigma_{i}/\Sigma_{p-i}\times\Sigma_{1}^{\times i}}\tau F_{p},(p-i,i)}(l,dl)=0$.
@@
Following Definition 7.3, the module of Kähler differentials $\Omega_{\Gamma(\operatorname{Lie})}(L)$ of $L$ is the $\mathbb{U}_{\Gamma(\operatorname{Lie})}(L)$-module generated by elements $dl$, $l\in L$, under the relations:
@@
$d([l,l^{\prime}])$ $=d(\beta_{[-,-],(1,1)}(l,l^{\prime}))=\beta_{[-,-],(1,1)}(-,l^{\prime})\otimes dl+\beta_{[-,-],(1,1)}(l,-)\otimes dl^{\prime}=[l,dl^{\prime}]-[l^{\prime},dl],$
@@
For all $i\in\{2,\dots,p-1\}$, note that again, $F_{p}=i\frac{1}{i}F_{p}=\sum_{\sigma\in\Sigma_{p-i}\times\Sigma_{i}/\Sigma_{p-i}\times\Sigma_{1}\times\Sigma_{i-1}}\frac{1}{i}F_{p}$. So, $\beta_{F_{p},(p-i,i)}(l,-)=\beta_{\sum_{\sigma\in\Sigma_{p-i}\times\Sigma_{i}/\Sigma_{p-i}\times\Sigma_{1}\times\Sigma_{i-1}}\frac{1}{i}F_{p},(p-i,i)}(l,-)$ is one of the elements described in Definition 6.1 in point (7), and so, is equal to 0 in $\mathbb{U}_{\Gamma(\operatorname{Lie})}(L)$. Finally, this last relation reads:
@@ 11.   Comparisons
The first part is proved for instance in [Rezk96]*Proposition 2.3.5. The second part follows from the first part and [Borceux94v2]*Proposition 4.3.2. ∎
@@
A regular epimorphism $q\colon X\twoheadrightarrow Y$ is the coequalizer of its kernel pair $X\times_{Y}X\rightrightarrows X$, which is a reflexive pair, with common section the diagonal $X\to X\times_{Y}X$. By 11.1, the functor $U^{\mathcal{P}}_{\mathbb{F}}$ preserves and reflects reflexive coequalizers, hence also regular epimorphisms.
@@
Since $\Gamma(\mathcal{P})\text{-}\mathrm{Alg}$ and $\mathcal{P}\text{-}\mathrm{Alg}$ are algebraic categories, 11.2 ensures that the adjunction
@@
(1) Via the explicit description of $\Gamma(\mathcal{P})$-algebras in 3.3, the restriction functor $f^{*}\colon\Gamma(\mathcal{Q})\text{-}\mathrm{Alg}\to\Gamma(P)\text{-}\mathrm{Alg}$ can be described as follows. Consider a $\Gamma(\mathcal{Q})$-algebra $A$ with operations $\beta_{y,\underline{r}}\colon A^{\times s}\to A$, given for all $\underline{r}=(r_{1},\dots,r_{s})$ and $y\in\mathcal{Q}(n)^{\Sigma_{\underline{r}}}$ with $n=r_{1}+\cdots+r_{s}$. Its restriction $f^{*}A$ has the same underlying $\mathbb{F}$-vector space $A$, with $\Gamma(\mathcal{P})$-algebra structure given by the operations
@@
(2) As observed in the proof of 11.2, in all four categories, regular epimorphisms are preserved and reflected by the forgetful functor to $\mathrm{Vect}_{\mathbb{F}}$. More concretely, they are the morphisms whose underlying map of vector spaces is surjective. ∎
Next, we want to compare Quillen cohomology of $\Gamma(\mathcal{P})$-algebras and $\mathcal{P}$-algebras. Start with a $\Gamma(\mathcal{P})$-algebra $A$ and consider its underlying $\mathcal{P}$-algebra $U_{\mathcal{P}}^{\Gamma(\mathcal{P})}A$, also denoted $A$ when the context indicates the category. The diagram of adjunctions in [Frankland15]*§4.2.2 specializes to
@@
Using the identification of Beck modules and Kähler differentials in [LodayV12]*Propositions 12.3.8 and 12.3.13 for the top row and 6.11 and 7.6 for the bottom row, the diagram becomes
@@
$\mu(a_{1},\ldots,a_{i-1},-,a_{i+1}\ldots,a_{n})\cdot\nu(g(b_{1}),\ldots,g(b_{j-1}),-,g(b_{j+1}),\ldots,g(b_{k}))\otimes_{\mathbb{U}_{\mathcal{P}}B}db_{j}$
@@
Here the left adjoint induced on Beck modules $\theta_{!}\colon\mathbb{U}_{\mathcal{P}}A\text{-}\mathrm{Mod}\to\mathbb{U}_{\Gamma(\mathcal{P})}A\text{-}\mathrm{Mod}$ is described in terms of the ring homomorphism $\theta$, but one might hope to describe it in terms of the original left adjoint $F^{\Gamma(\mathcal{P})}_{\mathcal{P}}\colon\mathcal{P}\text{-}\mathrm{Alg}\to\Gamma(\mathcal{P})\text{-}\mathrm{Alg}$. However, the left adjoint $F^{\Gamma(\mathcal{P})}_{\mathcal{P}}$ does not always pass to Beck modules in the sense of [Frankland15]*Definition 3.29. For example, in the case of divided power algebras, $F^{\Gamma(\mathrm{Com})}_{\mathrm{Com}}$ does not pass to Beck modules (see Proposition 12.3), while in the case of restricted Lie algebras, $F^{\Gamma(\operatorname{Lie})}_{\operatorname{Lie}}$ does pass to Beck modules (see Proposition 13.5).
@@ 12.   Divided power algebras versus commutative algebras
Let $A$ be a divided power algebra, equivalently, an object of $\mathrm{p}\mbox{-}\mathrm{Com}$. Since Beck modules over an augmented algebra $B$ are just $B$-modules, $A$ is the object $\mathbb{U}_{\mathrm{Com}}(A)$ representing Beck modules over the underlying augmented algebra of $A$. Recall from Section 9 that $V(A)=A\otimes_{\mathbb{F}}R_{f}$ represents Beck modules over $A$. The $\mathbb{F}$-linear ring homomorphism $\theta\colon A\to V(A)$ is given by $\theta(a)=a\otimes 1$. Denote by $\Omega^{1}_{A}$ the usual Kähler differentials over the underlying augmented algebra, whereas $\Omega_{\mathrm{p}\mbox{-}\mathrm{Com}}(A)$ was described in 9.7. The diagram of adjunctions (11.eq.2) specializes to
@@
Let By 9.5, a $V(A)$-module corresponds to a pair $(M,\pi)$, where $M$ is an $A$-module and $\pi\colon M\to M$ is a $p$-semilinear map such that $\pi(am)=0$ holds for all $a\in A_{+}$ and $m\in M$. Equivalently, the $V(A)$-module $(M,\pi)$ corresponds to an abelian group object $A\oplus_{p}{}M\to A$ in ($\mathrm{p}\mbox{-}\mathrm{Com}/A)_{\mathrm{ab}}$, where $A\oplus_{p}M$ is the semidirect product in $\mathrm{Com}$ of $A$ and $M$ together with the map
@@
Via 9.5, restriction of scalars along $\theta\colon A\to V(A)$ is the functor sending a pair $(M,\pi)$ to $M$, forgetting the $p$-semilinear map $\pi$.
@@
Part (1) follows from [Frankland15]*Proposition 4.12 and 12.1. Part (2) is a specialization of Proposition 11.5 (3) to the operad $\mathcal{P}=\mathrm{Com}$. ∎
@@
To conclude this section, we will specify the comparison maps of Proposition 11.6 to the case of divided power algebras. Let $g\colon B\to A$ be a morphism of non-unital divided power algebras. On the one hand,
@@ 13.   Restricted Lie algebras versus Lie algebras
Let $\mathbb{F}$ be a field of characteristic $p\neq 0$. Let $(H,(-)^{[p]})$ be a restricted Lie algebra. Then by 10.8 a $w(H)$-module $M$ is associated to a pair $({}_{u(H)}M,f)$, where ${}_{u(H)}M$ is $M$ viewed as a $u(H)$-module and $f\colon{}_{u(H)}M\to{}_{u(H)}M^{H}$ is a $p$-semilinear map. Equivalently, the $w(H)$-module $M$ is associated to the abelian group object
@@
    where $\theta\colon U(H)\to w(H)$ is the ring homomorphism described in 10.9.
@@
Part (1) follows from [Frankland15]*Proposition 4.12 and 13.2. Part (2) is a specialization of Proposition 11.5 (3) to the operad $\mathcal{P}=\operatorname{Lie}$. ∎
@@
From 13.4, we deduce that $g$ is in fact an isomorphism, hence $F^{\Gamma(\operatorname{Lie})}_{\operatorname{Lie}}(L\ltimes M\to L)$ has the structure of a Beck module over $F^{\Gamma(\operatorname{Lie})}_{\operatorname{Lie}}(L)$. ∎

### v3

@@ 10.2.   Operadic point of view
Remark 10.13.
In characteristic $p=2$, an algebra over the Lie operad, also called operadic Lie algebra, satisfies the antisymmetry equation $[x,y]=-[y,x]$ but might not satisfy the stronger condition $[x,x]=0$ [Fresse00]*Example 1.1.13. However, Theorem 10.12 holds also in characteristic $p=2$, since a $\Gamma(\operatorname{Lie})$-algebra satisfies the condition $[x,x]=0$ [Fresse00]*Remark 1.2.9.
We now use the characterisation of $\Gamma(\operatorname{Lie})$-algebras using Theorem 3.3. Following [Ikonicoff23]*Example 6.6.c, if $L$ is a $\Gamma(\operatorname{Lie})$-algebra, the Lie bracket on $L$ is given by $\beta_{[-,-],(1,1)}$, and the $p$-map is given by $\beta_{F_{p},(p)}$, where $F_{p}\in\operatorname{Lie}(p)^{\Sigma_{p}}$ is the element:
@@
where $\sigma$ ranges over $\Sigma_{1}\times\Sigma_{p}$ in the sum, which can also be written
$\beta_{[-,-],(1,1)}(\beta_{1_{\mathcal{P}},(1)},\beta_{F,(p)})=\beta_{F_{p+1},(1,p)}.$
@@
By Definition 4.5, an $L$-module $M$ becomes a vector space $M$ equipped with a semilinear map $f\colon M\to M$ and an action of $L$ that we denote $[l,m]$ for $l\in L$, $m\in M$ such that $[l,f(m)]=0$. This last relation comes from the following computation:
$[l,m^{[p]}]$ $=\beta_{[-,-],(1,1)}(l,\beta_{F,(p)}(m))$
$=\beta_{\sum_{\sigma}\sigma[-,-]\circ_{1}\cdots\circ_{1}[-,-],(1,p)}(l,m).$
@@
Note that $L$ injects into $\mathbb{U}_{\Gamma(\operatorname{Lie})}(L)$, by identifying each element $l\in L$ to $\beta_{[-,-],(1,1)}(l,-)$. For an $L$-module (i.e., a $\mathbb{U}_{\Gamma(\operatorname{Lie})}(L)$-module) $M$, we will denote by $[l,m]:=\beta_{[-,-],(1,1)}(l,m)=\beta_{[-,-],(1,1)}(l,-)\cdot m$, and $f(m):=\beta_{F_{p},(p)}(m)=\beta_{F_{p},(p)}(-)\cdot m$ for $l\in L$, $m\in M$.
@@
By Proposition 7.2, a Beck derivation $d\colon L\to M$ is a linear map $d\colon L\to M$ satisfying
$\begin{cases}d\left([l,l^{\prime}]\right)=\beta_{[-,-],(1,1)}(l,dl^{\prime})+\beta_{[-,-],(1,1)}(dl,l^{\prime})=[l,dl^{\prime}]-[l^{\prime},dl]\\ d\left(l^{[p]}\right)=\sum_{i=1}^{p}\beta_{F_{p},(p-i,i)}(l,dl).\\ \end{cases}$
Since $F_{p}\in Lie(p)^{\Sigma_{p}}$, for all $i$ such that $12\leq i\leq p-$ 1,
@@
and by relation ($\beta$M4), $\beta_{\sum_{\tau\in\Sigma_{p-i}\times\Sigma_{i}/\Sigma_{p-i}\times\Sigma_{1}^{\times i}}\tau F_{p},(p-i,i)}(l,dl)=0$.
@@
By Definition 7.3, the module of Kähler differentials $\Omega_{\Gamma(\operatorname{Lie})}(L)$ of $L$ is the $\mathbb{U}_{\Gamma(\operatorname{Lie})}(L)$-module generated by elements $dl$, $l\in L$, under the relations:
@@
$d([l,l^{\prime}])$ $=d(\beta_{[-,-],(1,1)}(l,l^{\prime}))$
$=\beta_{[-,-],(1,1)}(-,l^{\prime})\otimes dl+\beta_{[-,-],(1,1)}(l,-)\otimes dl^{\prime}$
$=[l,dl^{\prime}]-[l^{\prime},dl],$
@@
For all $i\in\{2,\dots,p-1\}$, again we have
$F_{p}=i\frac{1}{i}F_{p}=\sum_{\sigma\in\Sigma_{p-i}\times\Sigma_{i}/\Sigma_{p-i}\times\Sigma_{1}\times\Sigma_{i-1}}\frac{1}{i}F_{p}.$
So, $\beta_{F_{p},(p-i,i)}(l,-)=\beta_{\sum_{\sigma\in\Sigma_{p-i}\times\Sigma_{i}/\Sigma_{p-i}\times\Sigma_{1}\times\Sigma_{i-1}}\frac{1}{i}F_{p},(p-i,i)}(l,-)$ is one of the elements described in Definition 6.1 in point (7), and so, is equal to 0 in $\mathbb{U}_{\Gamma(\operatorname{Lie})}(L)$. This last relation reads:
@@ 11.   Comparisons
The first part is proved in [Rezk96]*Proposition 2.3.5. The second part follows from the first part and [Borceux94v2]*Proposition 4.3.2. ∎
@@
A regular epimorphism $q\colon X\twoheadrightarrow Y$ is the coequalizer of its kernel pair $X\times_{Y}X\rightrightarrows X$, which is a reflexive pair, with common section the diagonal $X\to X\times_{Y}X$. By Lemma 11.1, the functor $U^{\mathcal{P}}_{\mathbb{F}}$ preserves and reflects reflexive coequalizers, hence also regular epimorphisms.
@@
Since $\Gamma(\mathcal{P})\text{-}\mathrm{Alg}$ and $\mathcal{P}\text{-}\mathrm{Alg}$ are algebraic categories, Lemma 11.2 ensures that the adjunction
@@
(1) Via the explicit description of $\Gamma(\mathcal{P})$-algebras given in Theorem 8.2, the restriction functor $f^{*}\colon\Gamma(\mathcal{Q})\text{-}\mathrm{Alg}\to\Gamma(P)\text{-}\mathrm{Alg}$ can be described as follows. Consider a $\Gamma(\mathcal{Q})$-algebra $A$ with operations $\beta_{y,\underline{r}}\colon A^{\times s}\to A$, given for all $\underline{r}=(r_{1},\dots,r_{s})$ and $y\in\mathcal{Q}(n)^{\Sigma_{\underline{r}}}$ with $n=r_{1}+\cdots+r_{s}$. Its restriction $f^{*}A$ has the same underlying $\mathbb{F}$-vector space $A$, with $\Gamma(\mathcal{P})$-algebra structure given by the operations
@@
(2) As observed in the proof of Lemma 11.2, in all four categories, regular epimorphisms are preserved and reflected by the forgetful functor to $\mathrm{Vect}_{\mathbb{F}}$. More concretely, they are the morphisms whose underlying map of vector spaces is surjective. ∎
Next, we want to compare the Quillen cohomology of $\Gamma(\mathcal{P})$-algebras and $\mathcal{P}$-algebras. Start with a $\Gamma(\mathcal{P})$-algebra $A$ and consider its underlying $\mathcal{P}$-algebra $U_{\mathcal{P}}^{\Gamma(\mathcal{P})}A$, also denoted $A$ when the context indicates the category. The diagram of adjunctions in [Frankland15]*§4.2.2 specializes to
@@
Using the identification of Beck modules and Kähler differentials in [LodayV12]*Propositions 12.3.8 and 12.3.13 for the top row and Theorem 6.11 and Proposition 7.6 for the bottom row, the diagram becomes
@@
$\mu\left(a_{1},\ldots,a_{i-1},-,a_{i+1}\ldots,a_{n}\right)\cdot\nu\left(g(b_{1}),\ldots,g(b_{j-1}),-,g(b_{j+1}),\ldots,g(b_{k})\right)\otimes_{\mathbb{U}_{\mathcal{P}}B}db_{j}$
@@
Here the left adjoint $\theta_{!}\colon\mathbb{U}_{\mathcal{P}}A\text{-}\mathrm{Mod}\to\mathbb{U}_{\Gamma(\mathcal{P})}A\text{-}\mathrm{Mod}$ induced on Beck modules is described in terms of the ring homomorphism $\theta$, but one might hope to describe it in terms of the original left adjoint $F^{\Gamma(\mathcal{P})}_{\mathcal{P}}\colon\mathcal{P}\text{-}\mathrm{Alg}\to\Gamma(\mathcal{P})\text{-}\mathrm{Alg}$. However, the left adjoint $F^{\Gamma(\mathcal{P})}_{\mathcal{P}}$ does not always pass to Beck modules in the sense of [Frankland15]*Definition 3.29. For example, in the case of divided power algebras, $F^{\Gamma(\mathrm{Com})}_{\mathrm{Com}}$ does not pass to Beck modules (see Proposition 12.3), while in the case of restricted Lie algebras, $F^{\Gamma(\operatorname{Lie})}_{\operatorname{Lie}}$ passes to Beck modules (see Proposition 13.5).
@@ 12.   Divided power algebras versus commutative algebras
Let $A$ be a divided power algebra, equivalently, an object of $\mathrm{p}\mbox{-}\mathrm{Com}$. Since Beck modules over an augmented algebra $B$ are just $B$-modules, $A$ is the object $\mathbb{U}_{\mathrm{Com}}(A)$ representing Beck modules over the underlying augmented algebra of $A$. Recall from Section 9 that $V(A)=A\otimes_{\mathbb{F}}R_{f}$ represents Beck modules over $A$. The $\mathbb{F}$-linear ring homomorphism $\theta\colon A\to V(A)$ is given by $\theta(a)=a\otimes 1$. Denote by $\Omega^{1}_{A}$ the usual Kähler differentials over the underlying augmented algebra, whereas $\Omega_{\mathrm{p}\mbox{-}\mathrm{Com}}(A)$ was described in Theorem 9.7. The diagram of adjunctions (11.eq.2) specializes to
@@
Let By Theorem 9.5, a $V(A)$-module corresponds to a pair $(M,\pi)$, where $M$ is an $A$-module and $\pi\colon M\to M$ is a $p$-semilinear map such that $\pi(am)=0$ holds for all $a\in A_{+}$ and $m\in M$. Equivalently, the $V(A)$-module $(M,\pi)$ corresponds to an abelian group object $A\oplus_{p}{}M\to A$ in ($\mathrm{p}\mbox{-}\mathrm{Com}/A)_{\mathrm{ab}}$, where $A\oplus_{p}M$ is the semidirect product in $\mathrm{Com}$ of $A$ and $M$ together with the map
@@
Via Theorem 9.5, restriction of scalars along $\theta\colon A\to V(A)$ is the functor sending a pair $(M,\pi)$ to $M$, forgetting the $p$-semilinear map $\pi$.
@@
Part (1) follows directly from [Frankland15]*Proposition 4.12 and Lemma 12.1. Part (2) is a specialization of Proposition 11.5 (3) to the operad $\mathcal{P}=\mathrm{Com}$. ∎
@@
To conclude this section, we will specialize the comparison maps of Proposition 11.6 to the case of divided power algebras. Let $g\colon B\to A$ be a morphism of non-unital divided power algebras. On the one hand,
@@ 13.   Restricted Lie algebras versus Lie algebras
Let $\mathbb{F}$ be a field of characteristic $p\neq 0$. Let $(H,(-)^{[p]})$ be a restricted Lie algebra. Then by Theorem 10.8 a $w(H)$-module $M$ is associated to a pair $({}_{u(H)}M,f)$, where ${}_{u(H)}M$ is $M$ viewed as a $u(H)$-module and $f\colon{}_{u(H)}M\to{}_{u(H)}M^{H}$ is a $p$-semilinear map. Equivalently, the $w(H)$-module $M$ is associated to the abelian group object
@@
    where $\theta\colon U(H)\to w(H)$ is the ring homomorphism described in Lemma 10.9.
@@
Part (1) follows directly from [Frankland15]*Proposition 4.12 and Lemma 13.2. Part (2) is a specialization of Proposition 11.5 (3) to the operad $\mathcal{P}=\operatorname{Lie}$. ∎
@@
From Remark 13.4, we deduce that $g$ is in fact an isomorphism, hence $F^{\Gamma(\operatorname{Lie})}_{\operatorname{Lie}}(L\ltimes M\to L)$ has the structure of a Beck module over $F^{\Gamma(\operatorname{Lie})}_{\operatorname{Lie}}(L)$. ∎